OpenStudy (anonymous):

The global competition for the world cup is broken up into six continental zones. The number of teams in the Scottish team;s zone is 17 less than the number of teams in the puerto rican team's zone. There are a total of 88 teams in both zones. How many teams are in each zone? The global competition for the world cup is broken up into six continental zones. The number of teams in the Scottish team;s zone is 17 less than the number of teams in the puerto rican team's zone. There are a total of 88 teams in both zones. How many teams are in each zone? @Mathematics

OpenStudy (anonymous):

Let the number of teams in the Scottish zone be x. Then, number of teams in the puerto rican zone = x + 17 x + x + 17 = 88 2x + 17 = 88 2x = 88 - 17 = 71 But we can't divide 71 by two(It will come as 35.5). Did you make any typing error?

OpenStudy (anonymous):

Nope so is it (71,35.5)

OpenStudy (anonymous):

No, then it will be 35.5, 52.5

OpenStudy (anonymous):

But I'm not sure, how can the number of teams come in decimals?

OpenStudy (anonymous):

just round down then i guess

OpenStudy (anonymous):

It might be 36, 52 or 35, 53

OpenStudy (anonymous):

35,52

OpenStudy (anonymous):

No, if you add it, you'll get only 87 teams(not 88 as mentioned in the question).

OpenStudy (anonymous):

M-E-D-A-L

OpenStudy (anonymous):

Here is the answer: If we let the Puerto Rican zone be X, then we have that the number of teams in the Scottish zone is given by 88 - x. But we also have that the number of teams in the Scottish zone is given by 2X - 17. Now we make equal: 88 - X = 2X - 17 From here we get Team in Puerto Rican zone is 35 and the Scottish is 53.